OpenStudy (anonymous):

What is the approximate volume of the cone? Use 3.14 for π.

OpenStudy (anonymous):

what cone?

OpenStudy (anonymous):

The formula is V=1/3(pi)r^2h

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so...umm?

OpenStudy (anonymous):

You need an actual picture of a cone or something, I need to know the measurements...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So @Noemi95...

OpenStudy (anonymous):

is it 2170?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Okay let's see recall that the formula is V=1/3 (pi) r^2 h

OpenStudy (anonymous):

Let's plug in the values we know, so V= (1/3)(3.14)(12^2)(8)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Now we work the formula out, (12^2 =144) So, V=1205.76

OpenStudy (anonymous):

mmhmm

OpenStudy (anonymous):

do you get it now? Or, why the mhmm?

OpenStudy (anonymous):

not really, some of it

OpenStudy (anonymous):

What is it that confuses you?

OpenStudy (anonymous):

A.1206 B. 2170 C.3260 D.6510 ??? that part

OpenStudy (anonymous):

The answer would be A.1206. The result we got was 1205.76 & if you round it you get 1206.

OpenStudy (anonymous):

ohhh thank you, im taking a big test so heres ur medal

OpenStudy (anonymous):

it isn't as hard as it seems! (: and thank you